a paddle boat can move at a speed of 10km in still water. the boat is paddled 4km downstream in a river in the same time it takes to go 2km upstream. what is the speed of the river?
Answer by mananth(16949)

Boat speed = 10 km/h
current speed = x km/h

speed against current 10 - x
speed with current 10 + x

Distance against current 2 miles
Distance with current 4 miles

t=d/r time against = time with current -
2 / ( 10 - x )= 4 / ( 10 + x )
2 ( 10 + x )= 4 ( 10 -x )
20 + 2 x = 40 - -4 x
4 x 2 x 40 + -20
6 x = 20
/ 6
x = 3.33 km/h current speed